OpenStudy (xoxolove456):

Simplify: 450+25x= -4x^2+1296 + 450-25x

OpenStudy (alisoliman):

4x^2+50x-1296=0

OpenStudy (sila1453):

x = -25/4-sqrt(5809)/4 x = sqrt(5809)/4-25/4

OpenStudy (alisoliman):

Silva he or she just needs to simplify this equation.

OpenStudy (xoxolove456):

I just have to simplify that

OpenStudy (xoxolove456):

Not solve